# Revisiting Randomization in Greedy Model Search

### v1

@@ A.3.1 Proofs for Degrees of Freedom Results
[…]
where $\Phi(z)$ is the CDF of a Gaussian distribution. Using the Mills’ ratio bound, we have $\Phi(z)\leq 1-e^{-z^{2}}$ for $z$ large enough, which would imply that $(2\Phi(z)-1)^{N}\leq(1-2e^{-z^{2}})^{N}$. Since the integrand on the right hand side of (96) is decreasing, we have, for $N$ large enough, that
$\begin{split}\min_{0\leq t\leq\log N}1-(2\Phi(\sqrt{t})-1)^{N}&\geq 1-(2\Phi(\sqrt{\log N})-1)^{N}\\ &\geq 1-(1-2/N)^{N}\\ &\geq\frac{1}{2}.\end{split}$ (97)
Plugging this lower bound into (96) and integrating from 00 to $N$ completes the proof. ∎
Example A.1.
We wish to provide an example where RGS reduces training error through the mechanism of avoiding local optima that FS may fall into, as opposed to that of enlarging the hypothesis class. This means that we would like to compare FS to the base estimator of RGS (rather than the ensemble).
In the orthogonal feature setting, for the same value of $k$, the training loss of the base estimator is worse than FS, since we can do no better than selecting the $k$ features with the highest correlations with the response. Hence, we make use of a correlated feature setting. The idea, in a nutshell, is to create a regression vector that is strongly correlated with the response, but is not part of the optimal representation. Because of greediness, FS will select this spurious feature, but each base estimator of RGS may have a chance of avoiding its selection if it is omitted from the first candidate set $\mathcal{V}_{1}$.
More rigorously, consider a generative model with features $x_{1,p},x_{2,p},\ldots,x_{p,p}$. Also assume a feature $x_{0,p}$ which is unavailable for selection, but will be used to define the other features. Suppose that we have $\langle\mathbf{x}_{i,p},\mathbf{x}_{j,p}\rangle=\delta_{ij}$ for $0\leq i,j\leq{p-1}$ and that $x_{p,p}=3^{-1/2}(x_{1,p}+x_{2,p}+x_{0,p})$, so that $\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot p,p}\rangle=3^{-1/2}$ for $i=1,2$. Assume $y=\beta(x_{1,p}+x_{2,p})+\zeta\sum_{i=3}^{p-1}x_{i,p}$. In other words, we have noiseless observations. Assume further that $06^{-1/2}\beta>\zeta>$ 0.
Denote the features indices selected by FS (in order) as $\pi(1),\pi(2),\ldots,\pi(p)$. Since
$\langle\mathbf{x}_{\cdot i,p},\mathbf{y}\rangle=\begin{cases}\beta&i=1,2\\ \zeta&3\leq i\leq p-1\\ 2\cdot 3^{-1/2}\beta&i=p,\end{cases}$ (98)
@@
$\langle\mathbf{x}_{\cdot i,p},\mathbf{r}_{1}\rangle=\begin{cases}\beta/3&i=1,2\\ \zeta&3\leq i\leq p-1.\end{cases}$ (99)
@@
$\begin{split}\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|^{2}&=1-\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot p,p}\rangle^{2}\\ &=\begin{cases}2/3&i=1,2\\ 1&3\leq i\leq p-1.\end{cases}\end{split}$ (100)
@@
$\frac{|\langle\mathbf{x}_{\cdot i,p},\mathbf{r}_{1}\rangle|}{\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|}=\begin{cases}6^{-1/2}\beta&i=1,2\\ \zeta&3\leq i\leq p-1.\end{cases}$ (101)
@@
$\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}=(p-k)\zeta^{2}.$ (102)
We now study the base estimator of RGS, denoted $\check{f}_{k,m,p}$, which depends on a random seed $\theta$. At iteration $l$, suppose $p\notin\mathcal{M}_{l}$. Observe that the residual $r_{l}$ satisfies
$\frac{|\langle\mathbf{x}_{\cdot p,p},\mathbf{r}_{l}\rangle|}{\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|}=\begin{cases}2\cdot 3^{-1/2}\beta&\text{if}~{}|\{1,2\}\cap\mathcal{M}_{l}|=0\\ 2^{-1/2}\beta&\text{if}~{}|\{1,2\}\cap\mathcal{M}_{l}|=1\\ 0&\text{if}~{}|\{1,2\}\cap\mathcal{M}_{l}|=2.\end{cases}$ (103)
Hence, if both $x_{1,p}$ and $x_{2,p}$ are selected before $x_{p,p}$, then $x_{p,p}$ is never selected. Denote the feature indices selected by RGS (in order) as $\tau(1),\tau(2),\ldots,\tau(p)$. Then $\tau^{-1}(j)$ denotes the step at which feature $x_{j,p}$ was selected. Note that these depend on the random seed $\theta$. We now decompose the probability space into three events. Let
$\mathcal{E}_{1}=\{\tau^{-1}(1),\tau^{-1}(2)\leq\min(\tau^{-1}(p),k)\},$ (104)
$\mathcal{E}_{2}=\{\tau^{-1}(1),\tau^{-1}(2)\leq k\}\backslash\mathcal{E}_{1},$ (105)
$\mathcal{E}_{3}=\left(\mathcal{E}_{1}\cup\mathcal{E}_{2}\right)^{c}.$ (106)
@@
$\|\mathbf{y}-\check{\mathbf{f}}_{k,m,p}\|^{2}=\begin{cases}(p-k-1)\zeta^{2}&\text{if}~{}\mathcal{E}_{1}~{}\text{holds}\\ (p-k)\zeta^{2}&\text{if}~{}\mathcal{E}_{2}~{}\text{holds}\\ (p-k-3)\zeta^{2}+C(\theta)&\text{if}~{}\mathcal{E}_{3}~{}\text{holds},\end{cases}$ (107)
where $|C(\theta)|\leq 2\beta^{2}$. We next compute bounds for the event probabilities. For $l=1,\ldots,k$, let $S_{l}$ denote the set of $m$ indices sampled at iteration $l$. Denote $\gamma=m/p$ as before. By our previous discussion, we have
$\begin{split}\mathbb{P}(\mathcal{E}_{1})&\geq\mathbb{P}(1\in S_{1},0\notin S_{1},2\in S_{2},0\notin S_{2})\\ &=\frac{\binom{p-2}{m-1}\binom{p-3}{m-1}}{\binom{p}{m}\binom{p-1}{m}}\\ &=\gamma^{2}(1-\gamma^{2})+O(1/p).\end{split}$ (108)
@@
$\begin{split}\mathbb{P}(\mathcal{E}_{3})&=1-\mathbb{P}(\tau^{-1}(1),\tau^{-1}(2)\leq k)\\ &=\mathbb{P}(\tau^{-1}(1)>k~{}\text{or}~{}\tau^{-1}(2)>k)\\ &\leq\mathbb{P}(\tau^{-1}(1)>k)+\mathbb{P}(\tau^{-1}(2)>k)\\ &\leq\mathbb{P}\left(\sum_{l=1}^{k}W_{l}^{(1)}<3\right)+\mathbb{P}\left(\sum_{l=1}^{k}W_{l}^{(2)}<3\right)\\ &\leq 2e^{-k\gamma}(ek\gamma/2)^{2},\end{split}$ (109)
@@
$\begin{split}\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}&\geq\zeta^{2}\mathbb{P}(\mathcal{E}_{1})-2\beta^{2}\mathbb{P}(\mathcal{E}_{3})\\ &\geq\zeta^{2}(1-\gamma^{2})\gamma^{2}-4\beta^{2}e^{-k\gamma}(ek\gamma/2)^{2}-O(1/p).\end{split}$ (110)
@@
Appendix B Additional Simulations and Data
Here, we describe additional simulation results not included in the main paper
B.1 Prediction Performance
  • Banded correlation. In Figure 5 and Figure 6 respectively, we present the RISE and RTE results for exact sparsity, banded correlation with $\rho=0.5$, and aspect ratios $p/n=0.2,0.4,0.5,0.8$. As observed in the main paper, the increase in aspect ratio seems to slightly magnify the improvement of RGS over other methods.
  • Block correlation. In Figure 8 and Figure 8 respectively, we present the RISE and RTE results for exact sparsity, block correlation, and aspect ratio $p/n=0.1$. We set $\Sigma_{ij}=0.25$ if features $i\neq j$ lie in the same block, and 00 otherwise, with blocks comprising sets of 20 adjacent features. For this setting, RGS does not perform as well as in the banded settings at low SNR. We speculate that this is because the moderately correlated variables that live in the same block are almost indistinguishable when the signal is weak. At low SNR, RGS tends to pick several near-duplicates instead of one strong predictor, which inflates the effective degrees of freedom and mitigates the variance-reduction benefit we observe in the banded experiments.
  • Inexact sparsity. In Figure 9 and Figure 10 respectively, we present the RISE and RTE results for inexact sparsity, under both banded ($\rho=0.5$) and block correlation, aspect ratio $p/n=0.1$. Note that these results mirror the exact sparsity setting quite closely.
  • Varying correlation. In Figure 11 and Figure 12 respectively, we present the RISE and RTE results for exact sparsity, banded correlation with aspect ratios $p/n=0.1$, but varying the correlation strength $\rho\in\{0.1,0.25,0.5,0.75\}$. The improvement of RGS over other methods seems to be relatively robust to changes in correlation strength. However, at $\rho=0.75$ GS and its ensembles no longer vastly outperform lasso and elastic net at high SNR.
  • Heavy-tailed noise. In Figure 13 and Figure 14 respectively, we present the RISE and RTE results for exact sparsity, banded correlation with $\rho=0.5$ and block correlation, with aspect ratio $p/n=0.1$. However, instead of Gaussian noise, we use Laplace noise. The results are similar to the Gaussian noise setting.
  • Model misspecification. In Figure 15 and Figure 16, we present the results for RISE under a misspecified (nonlinear) model setting. We generate responses via $y_{i}=\eta g(\mathbf{x}_{i})+(1-\eta)\langle\mathbf{x}_{i},\boldsymbol{\beta}\rangle+\varepsilon_{i}$, where $g(\mathbf{x}_{i})$ is a nonlinear function comprising the squared terms and pairwise interactions for the first $5$ variables, scaled by $1/\sqrt{p}$. Here, $\eta\in[0,1]$ is a parameter that controls the degree of nonlinearity and is set at $\eta=0.5$ in Figure 15 and at $\eta=0.9$ in Figure 16. $\boldsymbol{\beta}$ follows the exact sparsity design ($\beta_{i}=1$ for $1\leq i\leq 10$ and $\beta_{i}=0$ otherwise). SNR is computed via the formula
    $\text{SNR}=\frac{1}{n\sigma^{2}}\sum_{i=1}^{n}\left[\eta\left(\sum_{j=1}^{5}x_{ij}^{2}+\sum_{j=1}^{4}\sum_{l=j+1}^{5}x_{ij}x_{il}\right)+(1-\eta)\sum_{j=1}^{10}x_{ij}\right]^{2},$ (111)
    and is varied over the grid $\{0.031,0.053,0.11,0.25,1.0\}$. We consider both banded correlation with $\rho=0.5$ and block correlation, with aspect ratio $p/n=0.1$. We see that RGS performs better than GS and competitively with all other methods across all settings.
(a) $p/n=0.2$
(b) $p/n=0.4$
(c) $p/n=0.5$
(d) $p/n=0.8$
Figure 5: The relative in-sample error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for banded correlated exact sparsity data.
(a) $p/n=0.2$
(b) $p/n=0.4$
(c) $p/n=0.5$
(d) $p/n=0.8$
Figure 6: The relative test error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for banded correlated exact sparsity data.
Figure 7: The relative in-sample error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for block correlated exact sparsity data.
Figure 8: The relative test error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for block correlated exact sparsity data.
(a) Banded
(b) Block
Figure 9: The relative in-sample error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for inexact sparsity data.
(a) Banded
(b) Block
Figure 10: The relative test error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for inexact sparsity data.
(a) $\rho=0.1$
(b) $\rho=0.5$
(c) $\rho=0.25$
(d) $\rho=0.75$
Figure 11: The relative in-sample error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for banded exact sparsity data over various correlation strengths.
(a) $\rho=0.1$
(b) $\rho=0.5$
(c) $\rho=0.25$
(d) $\rho=0.75$
Figure 12: The relative test error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for banded exact sparsity data over various correlation strengths.
(a) Banded
(b) Block
Figure 13: The relative in-sample error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for Laplacian noise
(a) Banded
(b) Block
Figure 14: The relative test error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for Laplacian noise
(a) Banded
(b) Block
Figure 15: The relative in-sample error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for model misspecification with $\eta=0.5$
(a) Banded
(b) Block
Figure 16: The relative in-sample error across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for model misspecification with $\eta=0.9$
B.2 Bias-Variance Tradeoff
In Figures 17, 18, 19, 20 and 21 respectively, we plot degrees of freedom vs. training MSE for RGS and GS at SNR values $0.031,0.053,0.11,0.5,1.0$. Each figure contains 4 panels, showing the results for (a) banded correlation ($\rho=0.5$) and exact sparsity, (b) banded correlation and inexact sparsity, (c) block correlation and exact sparsity, (d) block correlation and inexact sparsity. These results suggest that the tradeoff curve for RGS is generally shifted to the bottom left (relative to that of GS), with the shift being more significant at lower SNR settings and for banded correlation. For block correlation, RGS may even have larger degrees of freedom than GS for the same $k$, which is in accordance with its poor relative performance in Figure 8 and Figure 8.
(a) Banded Exact
(b) Banded Inexact
(c) Block Exact
(d) Block Inexact
Figure 17: The degrees of freedom for SNR $=0.031$ for various correlation and sparsity settings.
(a) Banded Exact
(b) Banded Inexact
(c) Block Exact
(d) Block Inexact
Figure 18: The Degrees of Freedom for SNR $=0.053$ for various correlation and sparsity settings.
(a) Banded Exact
(b) Banded Inexact
(c) Block Exact
(d) Block Inexact
Figure 19: The degrees of freedom for SNR $=0.11$ for various correlation and sparsity settings.
(a) Banded Exact
(b) Banded Inexact
(c) Block Exact
(d) Block Inexact
Figure 20: The Degrees of Freedom for SNR $=0.25$ for various correlation and sparsity settings.
(a) Banded Exact
(b) Banded Inexact
(c) Block Exact
(d) Block Inexact
Figure 21: The degrees of freedom for SNR $=1.0$ for various correlation and sparsity settings.

### v3

@@ Acknowledgements
$\begin{split}\mathbb{E}\left[\max_{1\leq i\leq N}Z_{i}\right]&=\int_{0}^{\infty}\mathbb{P}\left(\max_{1\leq i\leq N}Z_{i}\geq t\right)dt\\ &=\int_{0}^{\infty}1-\mathbb{P}\left(\max_{1\leq i\leq N}Z_{i}<t\right)dt\\ &=\int_{0}^{\infty}1-\mathbb{P}\left(Z_{1}<t\right)^{N}dt\\ &=\int_{0}^{\infty}1-(2\Phi(\sqrt{t})-1)^{N}dt,\end{split}$ (S.73)
where $\Phi(z)$ is the CDF of a Gaussian distribution. Using the Mills’ ratio bound, we have $\Phi(z)\leq 1-e^{-z^{2}}$ for $z$ large enough, which would imply that $(2\Phi(z)-1)^{N}\leq(1-2e^{-z^{2}})^{N}$. Since the integrand on the right hand side of (S.73) is decreasing, we have, for $N$ large enough, that
$\begin{split}\min_{0\leq t\leq\log N}1-(2\Phi(\sqrt{t})-1)^{N}&\geq 1-(2\Phi(\sqrt{\log N})-1)^{N}\\ &\geq 1-(1-2/N)^{N}\\ &\geq\frac{1}{2}.\end{split}$ (S.74)
Plugging this lower bound into (S.73) and integrating from 0 to $N$ completes the proof. ∎
S.1.5 Proof of Proposition 5.1
According to the decomposition (6), it suffices to prove
$\sum_{j=1}^{p}\hat{\beta}^{2}_{j,p}\big(w_{j}^{k,m,p}\big)^{2}\leq\sum_{j=1}^{p}\hat{\beta}^{2}_{j,p}\big(w_{j}^{k,m^{\prime},p}\big)^{2}.$ (S.75)
This follows immediately from Theorem A.3 in Chapter 5 from marshall11, where, in their notation, we take $g_{i}(z)=\beta^{2}_{i,p}z^{2}$.
S.1.6 Proof of Proposition 6.1
Denote the features indices selected by FS($p$) (in order) as $\pi(1),\pi(2),\ldots,\pi(p)$. Since
$\langle\mathbf{x}_{\cdot i,p},\mathbf{y}\rangle=\begin{cases}\beta&i=1,2\\ \zeta&3\leq i\leq p-1\\ 2\cdot 3^{-1/2}\beta&i=p,\end{cases}$ (S.76)
@@
$\langle\mathbf{x}_{\cdot i,p},\mathbf{r}_{1}\rangle=\begin{cases}\beta/3&i=1,2\\ \zeta&3\leq i\leq p-1.\end{cases}$ (S.77)
@@
$\begin{split}\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|^{2}&=1-\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot p,p}\rangle^{2}\\ &=\begin{cases}2/3&i=1,2\\ 1&3\leq i\leq p-1.\end{cases}\end{split}$ (S.78)
@@
$\frac{|\langle\mathbf{x}_{\cdot i,p},\mathbf{r}_{1}\rangle|}{\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|}=\begin{cases}6^{-1/2}\beta&i=1,2\\ \zeta&3\leq i\leq p-1.\end{cases}$ (S.79)
@@
$\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}=(p-k)\zeta^{2}.$ (S.80)
We now study the base estimator of EFS, denoted $\check{f}_{k,m,p}$, which depends on a random seed $\theta$. At iteration $l$, suppose $p\notin\mathcal{M}_{l}$. Observe that the residual $r_{l}$ satisfies
$\frac{|\langle\mathbf{x}_{\cdot p,p},\mathbf{r}_{l}\rangle|}{\|\mathbf{P}_{\mathcal{M}_{1}}^{\perp}\mathbf{x}_{\cdot i,p}\|}=\begin{cases}2\cdot 3^{-1/2}\beta&\text{if}~|\{1,2\}\cap\mathcal{M}_{l}|=0\\ 2^{-1/2}\beta&\text{if}~|\{1,2\}\cap\mathcal{M}_{l}|=1\\ 0&\text{if}~|\{1,2\}\cap\mathcal{M}_{l}|=2.\end{cases}$ (S.81)
Hence, if both $x_{1,p}$ and $x_{2,p}$ are selected before $x_{p,p}$, then $x_{p,p}$ is never selected. Denote the feature indices selected by EFS (in order) as $\tau(1),\tau(2),\ldots,\tau(p)$. Then $\tau^{-1}(j)$ denotes the step at which feature $x_{j,p}$ was selected. Note that these depend on the random seed $\theta$. We now decompose the probability space into three events. Let
$\mathcal{E}_{1}=\{\tau^{-1}(1),\tau^{-1}(2)\leq\min(\tau^{-1}(p),k)\},$ (S.82)
$\mathcal{E}_{2}=\{\tau^{-1}(1),\tau^{-1}(2)\leq k\}\backslash\mathcal{E}_{1},$ (S.83)
$\mathcal{E}_{3}=\left(\mathcal{E}_{1}\cup\mathcal{E}_{2}\right)^{c}.$ (S.84)
@@
$\|\mathbf{y}-\check{\mathbf{f}}_{k,m,p}\|^{2}\leq\begin{cases}(p-k-1)\zeta^{2}&\text{if}~\mathcal{E}_{1}~\text{holds},\\ (p-k)\zeta^{2}&\text{if}~\mathcal{E}_{2}~\text{holds},\\ (p-k)\zeta^{2}+2\beta^{2}&\text{if}~\mathcal{E}_{3}~\text{holds}.\end{cases}$ (S.85)
We next compute bounds for the event probabilities. For $l=1,\ldots,k$, let $S_{l}$ denote the set of $m$ indices sampled at iteration $l$. Denote $\gamma=m/p$ as before. By our previous discussion, we have
$\begin{split}\mathbb{P}(\mathcal{E}_{1})&\geq\mathbb{P}(1\in S_{1},0\notin S_{1},2\in S_{2},0\notin S_{2})\\ &=\frac{\binom{p-2}{m-1}\binom{p-3}{m-1}}{\binom{p}{m}\binom{p-1}{m}}\\ &=\gamma^{2}(1-\gamma^{2})+O(1/p).\end{split}$ (S.86)
@@
$\begin{split}\mathbb{P}(\mathcal{E}_{3})&=1-\mathbb{P}(\tau^{-1}(1),\tau^{-1}(2)\leq k)\\ &=\mathbb{P}(\tau^{-1}(1)>k~\text{or}~\tau^{-1}(2)>k)\\ &\leq\mathbb{P}(\tau^{-1}(1)>k)+\mathbb{P}(\tau^{-1}(2)>k)\\ &\leq\mathbb{P}\left(\sum_{l=1}^{k}W_{l}^{(1)}<3\right)+\mathbb{P}\left(\sum_{l=1}^{k}W_{l}^{(2)}<3\right)\\ &\leq 2e^{-k\gamma}(ek\gamma/2)^{2},\end{split}$ (S.87)
@@
$\begin{split}\mathbb{E}\big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}\big]-\mathbb{E}\big[\|\mathbf{y}-\check{\mathbf{f}}_{k,m,p}\|^{2}\big]&\geq\zeta^{2}\mathbb{P}(\mathcal{E}_{1})-2\beta^{2}\mathbb{P}(\mathcal{E}_{3})\\ &\geq\zeta^{2}(1-\gamma^{2})\gamma^{2}-4\beta^{2}e^{-k\gamma}(ek\gamma/2)^{2}-O(1/p).\end{split}$ (S.88)